A submarine is on the surface of the ocean. The submarine then dives below the surface and travel a distance of 450 m while dropping a vertical distance of 300 m. determine the angle of depression and how far the submarine traveled horizontally
1
Expert's answer
2019-10-18T09:38:33-0400
"d^2=x^2+y^2"

Given that "d=450m,y=300m."


"x=\sqrt{d^2-y^2}"

"x=\sqrt{(450m)^2-(300m)^2}=150\sqrt{5}m\approx335.41m"

Determine the angle of depression


"\sin\theta={y\over d}"

"\theta=\arcsin{y\over d}"

"\theta=\arcsin{300m\over 450m}=\arcsin{2\over 3}\approx41.81\degree"

The angle of depression is "\theta=\arcsin{2\over 3}\approx41.81\degree."

The submarine traveled "150\sqrt{5}m\approx335.41m" horizontally.